Two positive charges of magnitude 2.4 x 10-9 C and 1.6 x 10-9 C are separated in air by a distance of 2cm. how much electrostatic potential is produced by the two charges at midway between them? 


1
Expert's answer
2021-05-30T13:32:11-0400

Find the potential from each charge at the midway:


"V_1=\frac{kq_1}{r}=2157\text{ V}.\\\space\\
V_2=\frac{kq_2}{r}=1438\text{ V}.\\\space\\
V=V_1+V_2=3595\text{ V}."